A uniform metre rule balanes on a knife edge placed at the 60cm mark whem a weight of 2N is suspended from one end calculate the weight of the metre rule
1
Expert's answer
2020-05-25T11:04:37-0400

The ruler weights m kilograms.



(there is nothing interesting on the 60-cm end, that is why it is not shown)


Since the ruler is uniform, we can say that the centre of mass of the 60-cm long side is at the centre, or the lever 30 cm. The 60-cm long part weights


"P_{60}=mg\frac{60}{100}."

The centre of mass of the 40-cm end is at the centre as well, the lever for this is 20 cm. The weight is


"P_{40}=mg\frac{40}{100}."

On the 40-cm end, there is 2 N attached at the 100-cm mark. The lever for 2 N is 40 cm. The system is in equilibrium, which means


"\sum d_{i\text{ left}}P_{i\text{ left}}=\sum d_{i\text{ right}}P_{i\text{ right}}\\
30P_{60}=20P_{40}+40\cdot P_\text{2N},\\
30\cdot mg\frac{60}{100}=20\cdot mg\frac{40}{100}+40\cdot2,\\
m=0.82\text{ kg}."